# Spatial-Temporal Decoupling Contrastive Learning for Skeleton-based Human Action Recognition

### v3

@@ 4.4 Comparsion With Other Contrastive Learning Methods
We compare our method with the recent contrastive learning methods in supervised skeleton-based action recognition. The comparison is conducted on the X-sub benchmark of the NTU120 dataset. As shown in Table 3, our STD-CL achieves competitive improvements. Furthermore, most previous methods are only compatible with GCN-based methods. To take advantage of our method, we combine the STD-CL with Hyperformer Zhou et al. (2022), which is a transformer-based method. Our method outperforms previous contrastive learning methods by 1.2%, which is a promising result. For more comparisons, please refer to the supplementary materials.
4.5 Ablation Study
Figure 5: Visualization of feature representation by t-SNE from the sequences in the test set of the X-sub benchmark of NTU120 dataset. Different colors denote different classes. The left is from the CTR-GCN, while the right is from our method.
In this section, we evaluate the different experimental settings on the X-sub benchmark of the NTU120 dataset to verify the design of the proposed STD-CL. For more ablation studies, please refer to the supplementary materials.
Contrastive Strategies. In Table 3, we conduct experiments to validate the contrastive strategy. We employ the decoupled spatial features or decoupled temporal features to contrastive learning, and these lead to performance improvements, which are owed to the decrease in the degree of mixture in the global features. Furthermore, we find that global feature contrast can improve the performance on the baseline with 0.5% (84.9 % to 85.6 %), which owes to the explicit modeling of latent data distributions. Furthermore, we combine the global feature contrast and our STD-CL, and the performance drops slightly (85.6% to 85.5%), which may be due to that the global features contrast limits the feature decoupling to some extent.
Table 4: Ablation study about the contrastive strategies. SD-CL means applying spatial decoupled features to contrastive learning, and TD-CL means applying temporal decoupled features to contrastive learning. G-CL means applying the global features to contrastive learning.
@@ 4.5 Ablation Study
Baseline (w/o contrast) 84.9
w/SD-CL $\text{85.4}^{\uparrow{0.5}}$
w/TD-CL $\text{85.3}^{\uparrow{0.4}}$
w/G-CL $\text{85.4}^{\uparrow{0.5}}$
w/STD-CL + w/ G-CL $\text{85.5}^{\uparrow{0.6}}$
w/STD-CL $\text{85.6}^{\uparrow{0.7}}$
Performance on Different Level Actions. To evaluate the effectiveness of the proposed STD-CL, we conduct recognition accuracy for different action categories with different levels of difficulty. Specifically, we gather actions whose accuracy is over 90% as Easy Level, between 80% to 90% as Medium Level, and lower than 80% as Hard Level for their respective classification results. The experimental results are displayed in Table 5. From the results, we can observe that the proposed STD-CL achieves relatively significant improvements in the hard and medium-level actions, which include more fine-grained action categories. Moreover, the proposed STD-CL has made some gains in the easy-level actions, thanks to the explicit modeling for the latent data distributions of contrastive learning.
@@
2S-AGCN w/STD-CL $\text{65.9}^{\uparrow{0.9}}$ $\text{86.6}^{\uparrow{0.4}}$ $\text{95.2}^{\uparrow{0.1}}$
@@
CTR-AGCN w/STD-CL $\text{66.1}^{\uparrow{0.2}}$ $\text{87.0}^{\uparrow{1.2}}$ $\text{95.9}^{\uparrow{0.5}}$
@@
Hyperformer w/STD-CL $\text{69.6}^{\uparrow{1.3}}$ $\text{87.0}^{\uparrow{1.1}}$ $\text{95.4}^{\uparrow{0.2}}$
Furthermore, we compare our results with a SOTA model CTR-GCN Chen et al. (2021a), and the results of class with accuracy differences higher than 3% between CTR-GCN and our method are displayed in Figure 4. We can obverse that actions such as “hit with object”, “point to something”, “wipe face”, “reading” and “yawn” benefit from the proposed STD-CL due to the great capacity to obtain discriminative and semantically distinct features. Besides, the reason for the poor performance of actions “take object out of bag”, “put on a shoe”, and “tear up paper” is that such actions are strongly object-related, making it challenging to recognize and contrast.
Training Consumption. In Table 6, we report the training consumption on NTU120. With our method, the training memory usage only slightly increases with different baseline models, and the increase in training time remains within an acceptable range, which proves the efficiency of the design.
@@ 4.6 Qualitative Analysis
In this section, we validate our STD-CL through t-SNE Van der Maaten and Hinton (2008) distribution visualization of feature representations in the test set of the X-sub benchmark of the NTU120 dataset. As shown in Figure 5, our STD-CL can exploit the data distributions explicitly with more compact intra-class distances and clear classification boundaries. Thus, the features extracted from the proposed STD-CL are more discriminative, which shows that the decoupled feature contrast improves the feature extraction capacity of the skeleton encoders.
@@ 5 Conclusion
In this paper, we propose a novel spatial-temporal decoupling contrastive learning (STD-CL) framework for skeleton-based action recognition, which can be combined with most of the previous skeleton encoders. We employ contrastive learning for the spatial and temporal decoupled features to encourage the skeleton encoders to extract more discriminative and semantically distinct features, which are proven to be effective for recognizing fine-grained action classes and ambiguous samples. The extensive experimental results with four various skeleton encoders on three datasets demonstrate the effectiveness of the proposed framework.

### v4

@@ 4.5 Ablation Study
In this section, we evaluate the different experimental settings on the X-sub benchmark of the NTU120 dataset to verify the design of the proposed CSRE.
Contrastive Strategy. In Table 5, we conduct experiments to validate the contrastive strategy. We employ the decomposed spatial-aware features or decomposed temporal-aware features to contrastive learning, and these lead to performance improvements, which are owed to the exploration of fine-grained motion patterns along the corresponding dimension. Furthermore, we find that global feature contrast can improve the performance on the baseline with 0.5% (84.9 % to 85.6 %), which owes to the explicit modeling of latent data distributions. Furthermore, we combine the global feature contrast and our CSRE, and the performance drops slightly (85.6% to 85.5%), which may be due to that the global features contrast limits the feature decomposition to some extent.
Table 5: Ablation study about the contrastive strategies. SC means applying spatial-specific features to contrastive learning, and TC means applying temporal-specific features to contrastive learning. GC means applying the global feature to contrastive learning. Bold indicates the best results.
@@
CTR-GCN (w/o contrast) 84.9
CTR-GCN w/SC $\text{85.4}^{\uparrow{0.5}}$
CTR-GCN w/TC $\text{85.3}^{\uparrow{0.4}}$
CTR-GCN w/GC $\text{85.4}^{\uparrow{0.5}}$
CTR-GCN w/CSRE + w/GC $\text{85.5}^{\uparrow{0.6}}$
CTR-GCN w/CSRE $\textbf{85.6}^{\uparrow{0.7}}$
Sampling Strategy. In Table 6, we conduct experiments to validate the sampling strategy. In the training phase of the proposed CSRE, we adopt a hybrid sampling strategy to obtain samples from a memory bank. Specifically, hard mining of positive samples and the combination of random sampling and hard mining of negative samples.
Table 6: Ablation study about sampling strategy. R: Random; H: Hard mining. Bold indicates the best results.
Sampling Acc(%)
Positive Negative
R R 85.1
H 85.2
R + H 85.4
H R 85.2
H 85.3
R + H 85.6
R + H R + H 85.4
Performance on Different Level Actions. To evaluate the effectiveness of the proposed CSRE, we conduct recognition accuracy for different action categories with different levels of difficulty. Specifically, we gather actions whose accuracy is over 90% as Easy Level, between 80% to 90% as Medium Level, and lower than 80% as Hard Level for their respective classification results. The experimental results are displayed in Table 7. From the results, we can observe that the proposed CSRE achieves relatively significant improvements in the hard and medium-level actions, which include more fine-grained action categories. Moreover, the proposed CSRE has made some gains in the easy-level actions, which is possible thanks to the explicit modeling of the latent data distribution of contrastive learning.
@@
2S-AGCN w/CSRE $\text{65.9}^{\uparrow{0.9}}$ $\text{86.6}^{\uparrow{0.4}}$ $\text{95.2}^{\uparrow{0.1}}$
@@
CTR-AGCN w/CSRE $\text{66.1}^{\uparrow{0.2}}$ $\text{87.0}^{\uparrow{1.2}}$ $\text{95.9}^{\uparrow{0.5}}$
@@
Hyperformer w/CSRE $\text{69.6}^{\uparrow{1.3}}$ $\text{87.0}^{\uparrow{1.1}}$ $\text{95.4}^{\uparrow{0.2}}$
Furthermore, we compare our results with a SOTA model CTR-GCN [5], and the results of class with accuracy differences higher than 3% between CTR-GCN and our method are displayed in Figure 5. We can observe that actions such as "hitting with object", "pointing to something", "wipe face", "reading" and "yawn" benefit from the proposed CSRE due to the great capacity to obtain discriminative and semantically distinct features. Besides, the reason for the poor performance of actions "take object out of bag", "put on a shoe", and "tear up paper" is that such actions are strongly object-related, making it challenging to recognize and contrast.
Training Consumption. In Table 8, we report the training consumption on NTU120. With our method, the training memory usage only slightly increases with different baseline models, and the increase in training time remains within an acceptable range, which proves the efficiency of the design. Besides, although we set the two memory banks to store the features of all samples, as the size of the dataset increases, it is necessary to balance computational consumption and performance by limiting the capacity of memory banks. Similar to MoCo[28], using a queue-based mechanism to manage the memory bank capacity is a viable approach.
@@ 4.6 Qualitative Analysis
Figure 6: Visualization of feature representation by t-SNE from the sequences in the test set of the X-sub benchmark of NTU120 dataset. Different colors denote different classes.
In this section, we validate our CSRE through t-SNE [9] distribution visualization of feature representations in the test set of the X-sub benchmark of the NTU120 dataset. As shown in Figure 6, all contrastive learning methods, including proposed CSRE, learn more discriminative representations compared to the CTR-GCN [5]. Moreover, it can be observed that the feature distribution learned by CSRE has more significant classification boundaries than other contrastive learning methods with fewer ambiguous samples. Thus, the features extracted from the proposed CSRE are more discriminative, which shows that the CSRP improves the feature extraction capacity of the skeleton encoders.
@@ 5 Conclusion
In this paper, we propose a novel and generically contrastive spatiotemporal representation enhancement (CSRE) framework for skeleton-based action recognition, which can be combined with most previous skeleton encoders. We employ contrastive learning for the spatial-specific and temporal-specific features to encourage the skeleton encoders to enhance the distinctiveness of the representation, which is proven to be effective for recognizing fine-grained action classes and ambiguous samples. The extensive experimental results with five various skeleton encoders on five benchmarks demonstrate the effectiveness of the proposed framework.
Limitation and Discussion. Although the proposed CSRE performs solid improvements on challenging datasets, it also suffers from some dilemmas. As demonstrated in section 4.5, the performer with CSRE is declining in some specific action categories. Since the features cannot form stable clusters during training, CSRE may not have a positive improvement in some categories that originally performed poorly. We suppose this may be due to the poor performance caused by insufficient spatiotemporal modeling of the skeleton encoders. We will do further research on the above aspect in future work.
Acknowledgments
This work was supported partly by the National Natural Science Foundation of China (Grant No. 62173045), partly by the Natural Science Foundation of Hainan Province (Grant No. 622RC675), the Fundamental Research Funds for the Central Universities (Grant No. 2020XD-A04-3), and the 2023 “Double First-Class” Discipline Construction Interdisciplinary Team Project of Beijing University of Posts and Telecommunications (Grant No.2023SYLTD02).